sir, I have asked one problem sum in physics couple of days ago,i have not received the solution yet. kindly explain how to work out to find ½ Nb- Na,bonding and anti bonding for N2 and O2 in Molecuar orbital theory.

sir,
 
I have asked one problem sum in physics couple of days ago,i have not received the solution yet.
 
kindly explain how to work out to find  ½ Nb- Na,bonding and anti bonding for N2 and O2 in Molecuar orbital theory.
 
 

Grade:12

1 Answers

Mahima Kanawat
5 years ago
Dear student 
First of all you need to write configuration of N2 and O2....
Configuration of N2 (σ1s)2(σ∗1s)2(σ2s)2(σ∗2s)2(π2px)2(π 2py)2(σ2pz)2    
Here Nb = 10, Na = 4
So, Nb-Na/2 = 10-4/2=6/2=3
WITH REGARDS 
MAHIMA 
ASKIITIANS FORUM EXPERT 
Configuration of O2